# An Efficient Algorithm for Determining the Equivalence of Zero-one Reaction Networks

### v1

@@ Abstract
Zero-one reaction networks play a crucial role in cell signaling. Determining the equivalence of reaction networks is a fundamental computational problem in the field of chemical reaction networks. In this work, we develop an efficient method for determining the equivalence of zero-one networks. The efficiency comes from several criteria for determining the equivalence of the steady-state ideals arising from zero-one networks, which helps for cutting down the expenses on computing Gröbner bases. Experiments show that our method can successfully classify over three million networks according to their equivalence in a reasonable time.
@@ 1 Introduction
For the dynamical systems that arise from biochemical reaction networks, the following computational question is fundamental.
@@
The equivalence of steady-state ideals is one of the equivalence problem in the field of chemical reaction networks. The dynamical features of chemical reaction networks such as multistability, Hopf-bifurcation, and absolute concentration robustness (ACR) are linked to switch-like behavior, decision-making process, oscillation, and so on in cellular signaling [7, 23, 27, 9, 3, 25]. Usually, looking at the steady states is the first step for studying the dynamical behaviors of a chemical reaction system arising under mass-action kinetics. However, it is possible for two different networks to have the same steady-state ideal (notice that the steady states can be considered as real points located in the algebraic variety generated by the steady-state ideal). For instance, consider the two reaction networks:
$X_{1}+X_{2}\xrightarrow{\kappa_{1}}0,$ $X_{2}\xrightarrow{\kappa_{2}}X_{1}+X_{2},$ $0\xrightarrow{\kappa_{3}}X_{1}+X_{2},$ $0\xrightarrow{\kappa_{4}}X_{2};$ (1)
$X_{1}+X_{2}\xrightarrow{\kappa_{1}}X_{2},$ $X_{2}\xrightarrow{\kappa_{2}}X_{1},$ $0\xrightarrow{\kappa_{3}}X_{1},$ $0\xrightarrow{\kappa_{4}}X_{2}.$ (2)
They respectively have the following steady-state systems (see how to write down the system for a given network later in (8))
$\{-\kappa_{1}x_{1}x_{2}+\kappa_{2}x_{2}+\kappa_{3},-\kappa_{1}x_{1}x_{2}+\kappa_{3}+\kappa_{4}\};$ (3)
$\{-\kappa_{1}x_{1}x_{2}+\kappa_{2}x_{2}+\kappa_{3},-\kappa_{2}x_{2}+\kappa_{4}\}.$ (4)
It is straightforward to compute the reduced Gröbner bases [8] for the above two steady-state systems, and we find that in $\mathbb{Q}(\kappa)[x]$ they have the same reduced Gröbner basis
@@
In many cases, the above two networks are considered to be equivalent because when we only care about the dynamical behaviors determined by the steady states such as multistationarity, multistability, Hopf-bifurcation and ACR, we only need to study one of the networks that have the same steady-state ideal. Notice that relabeling the species or the reactions in a network naturally gives another kind of equivalence (see more details in [5] on the “dynamical equivalence” of networks). For instance, if we relabel the species $X_{1}$ and $X_{2}$ and we relabel the first and the second reactions in the network (1), then we get the following network
[…]
Of course, the network (6) should be considered to be equivalent to the network (1) and so, it is also equivalent to the network (2). Hence, in our context, we say two networks are equivalent if they have the same steady-state ideal after relabeling (see Definition 1).
Determining the equivalence is computationally challenging since a standard approach is to compute the reduced Gröbner bases. In a recent related work [12], the authors point out that there are few works on looking at the equivalence of the steady-state ideals, and we see that deriving explicit conditions for the equivalence is difficult in general. So, we might need to focus on a special class of networks for making this problem more feasible. The standard way (computing Gröbner bases [17, 15, 13, 14, 1, 16]) might be efficient for small networks such as one-species networks or two-reaction networks. However, in the studying of small networks, we usually need to deal with a substantial amount of networks. The idea of studying the small networks is motivated by the fact that many important dynamical features have inheritance from a large network to a smaller “subnetwork” such as multistability [20, 6], oscillation [2] and local bifurcations [4]. In many recent works, people are searching for the smallest networks from a big class of networks that admit these dynamical behaviors. In these studies, we need to first determine the equivalence of networks for making the searching process easier. For instance, Banaji and Boros [3] recently completely classify the smallest at-most-bimolecular networks that admit Hopf bifurcations (they contain three species and four reactions). Tang and Wang [27] find that the smallest zero-one networks that admit Hopf bifurcations are four-dimensional (they contain four species and five reactions). Kaihnsa, Nguyen and Shiu [21] prove that an at-most-bimolecular network admitting both multistationarity and ACR has at least three species and three reactions, and it is at least two-dimensional. The focus of this work is the zero-one networks since many biological significant networks in cell signaling are zero-one such as phosphorylation-dephosphorylation cycle [18], cell cycle [24], hybrid histidine kinase [26], and so on.
In this work, our main contribution is an efficient algorithm (Algorithm 1) for determining the equivalence of zero-one networks. The efficiency of the proposed new method comes from several criteria for determining the equivalence of the steady-state ideals arising from zero-one networks, which dramatically cuts down the expenses on computing Gröbner bases since many computations are avoided. We have successfully applied the new method to determining the equivalence of the smallest zero-one networks that admit multistationarity (these networks are three-dimensional, three-species and five-reaction [19]). The input contains over three millions such networks, and the new algorithm returns $32394$ networks after determining the equivalence. The computation is completed in a short time (see Table 1).
@@
The rest of this paper is organized as follows. In Section 2, we review the basic concepts on reaction networks and the dynamical systems arising from mass-action kinetics. Also, we formally define the steady-state ideal and the equivalence of networks in our context. In Section 3, we prove a series of sufficient conditions (Theorems 13) for determining the equivalence of networks. In Section 4, we present the main algorithm (Algorithm 1) with two sub-algorithms (Algorithms 23) for determining the equivalence of zero-one networks. In Section 5, we illustrate how the main algorithm performs by applying it to the three-dimensional zero-one networks with three species and five reactions. Also, we explain the implementation details and present the computational timings.
@@ 2 Background
where all stoichiometric coefficients $\alpha_{ij}$ and $\beta_{ij}$ are non-negative integers, and we assume that $(\alpha_{1j},\ldots,\alpha_{sj})\neq(\beta_{1j},\ldots,\beta_{sj})$. Each $\kappa_{j}\in\mathbb{R}_{>0}$ is called a rate constant corresponding to the $j$-th reaction in (7). We say a reaction is a zero-one reaction, if the stoichiometric coefficients $\alpha_{ij}$ and $\beta_{ij}$ in (7) belong to $\{0,1\}$. We say a network (7) is a zero-one network if it only has zero-one reactions. For any reaction defined in (7), the stoichiometric vector is $(\beta_{1j}-\alpha_{1j},\dots,\beta_{sj}-\alpha_{sj})^{\top}\in\mathbb{Q}^{s}$. We say finitely many reactions are linearly independent if their corresponding stoichiometric vectors are linearly independent. We call the $s\times m$ matrix with $(i,j)$-entry equal to $\beta_{ij}-\alpha_{ij}$ the stoichiometric matrix of $G$, denoted by $\mathcal{N}$. The dimension of $G$ is defined as the rank of $\mathcal{N}$. We call the $s\times m$ matrix the reactant matrix of $G$ with $(i,j)$-entry equal to $\alpha_{ij}$, denoted by $\mathcal{Y}$.
@@
$\dot{x}\;=\;f(\kappa,x)\;:=\;\mathcal{N}v(\kappa,x)\;=\;\mathcal{N}\begin{pmatrix}\kappa_{1}\prod\limits_{i=1}^{s}x_{i}^{\alpha_{i1}}\\ \kappa_{2}\prod\limits_{i=1}^{s}x_{i}^{\alpha_{i2}}\\ \vdots\\ \kappa_{m}\prod\limits_{i=1}^{s}x_{i}^{\alpha_{im}}\\ \end{pmatrix},$ (8)
@@
By viewing the rate constants as a real vector of parameters $\kappa:=(\kappa_{1},\kappa_{2},\dots,\kappa_{m})^{\top}$, we have polynomials $f_{i}(\kappa,x)\in\mathbb{Q}(\kappa)[x]$, for $i\in\{1,\dots,s\}$. The polynomial ideal generated by $f$ in $\mathbb{Q}(\kappa)[x]$, denoted by $\langle f\rangle$, is called the steady-state ideal.
@@
Two reaction networks are equivalent if they have the same steady-state ideal in the sense of relabeling.
For any given rate-constant vector $\kappa^{*}\in{\mathbb{R}}^{m}_{>0}$, a steady state of (8) is a vector of concentrations $x^{*}\in\mathbb{R}_{\geq 0}^{s}$ such that $f(\kappa^{*},x^{*})=\vec{0}$, where $f(\kappa,x)$ is on the right-hand side of the ODEs (8), and $\vec{0}$ denotes the column vector whose coordinates are all zero. If all the coordinates of a steady state $x^{*}$ are strictly positive (i.e., $x^{*}\in\mathbb{R}_{>0}^{s}$), then we call $x^{*}$ a positive steady state. For the stoichiometric matrix $\mathcal{N}\in\mathbb{Q}^{s\times m}$, the positive flux cone of $\mathcal{N}$ is defined as
@@
In applications, we usually only care about positive steady states of a network. It is well-known that a network admits positive steady states if and only if $\mathcal{F}^{+}(\mathcal{N})\neq\emptyset$.

### v2

# Determining the Equivalence of Small Zero-one Reaction Networks

@@ Abstract
Zero-one reaction networks are pivotal to cellular signaling, and establishing the equivalence of such networks represents a foundational computational challenge in the realm of chemical reaction network research. Herein, we propose a high-efficiency approach for identifying the equivalence of zero-one networks. Its efficiency stems from a set of criteria tailored to judge the equivalence of steady-state ideals derived from zero-one networks, which effectively reduces the computational cost associated with Gröbner basis calculations. Experimental results demonstrate that our proposed method can successfully categorize more than three million networks by their equivalence within a feasible timeframe. Also, our computational results for two important classes of quadratic zero-one networks ($3$-dimensional with $3$ species, $6$ reactions; $4$-dimensional with $4$ species, $5$ reactions) show that they have no positive steady states for a generic choice of rate constants, implying these small networks generically exhibit neither multistability nor periodic orbits.
@@ 1 Introduction
The dynamical systems generated by biochemical reaction networks give rise to the following fundamental computational question.
@@
The congruence of steady-state ideals is one of the core equivalence problems in the field of chemical reaction networks. Numerous key dynamical characteristics of chemical reaction networks—including multistability, Hopf bifurcations, and absolute concentration robustness (ACR)—are correlated with switch-like behaviors, decision-making processes, oscillations, and other key events in cellular signaling [8, 22, 26, 10, 3, 24]. Typically, examining the steady states constitutes the primary step in investigating the dynamical behaviors of a chemical reaction system arising under mass-action kinetics. Notice that the steady states can be considered as positive real points located in the algebraic variety generated by the steady-state ideal. However, it is possible for two distinct networks to have exactly the same steady-state ideal. For instance, consider the following two reaction networks (1) and (2).
$X_{1}+X_{2}\xrightarrow{\kappa_{1}}0,$ $X_{2}\xrightarrow{\kappa_{2}}X_{1}+X_{2},$ $0\xrightarrow{\kappa_{3}}X_{1}+X_{2},$ $0\xrightarrow{\kappa_{4}}X_{2}$ (1)
$X_{1}+X_{2}\xrightarrow{\kappa_{1}}X_{2},$ $X_{2}\xrightarrow{\kappa_{2}}X_{1},$ $0\xrightarrow{\kappa_{3}}X_{1},$ $0\xrightarrow{\kappa_{4}}X_{2}$ (2)
Their steady-state ideals are respectively generated by the following two sets of polynomials (3) and (4) (we will explain how to write down these polynomials for the given networks later in (8)).
$\{-\kappa_{1}x_{1}x_{2}+\kappa_{2}x_{2}+\kappa_{3},-\kappa_{1}x_{1}x_{2}+\kappa_{3}+\kappa_{4}\}$ (3)
$\{-\kappa_{1}x_{1}x_{2}+\kappa_{2}x_{2}+\kappa_{3},-\kappa_{2}x_{2}+\kappa_{4}\}$ (4)
It is straightforward to compute that the above two steady-state systems have the same reduced Gröbner basis [9] in $\mathbb{Q}(\kappa)[x]$:
@@
In numerous instances, the two networks above are regarded equivalent if our focus is solely on dynamical behaviors governed by steady states, including multistationarity, multistability, Hopf bifurcations, and ACR. Notice that relabeling the species or the reactions in a network naturally gives another kind of equivalence, say natural equivalence. For example, if we relabel the species $X_{1}$ and $X_{2}$ and we relabel the first and the second reactions in network (1), then we get the following network
$X_{1}\xrightarrow{\kappa_{1}}X_{1}+X_{2},\;\;X_{1}+X_{2}\xrightarrow{\kappa_{2}}0,\;\;0\xrightarrow{\kappa_{3}}X_{1}+X_{2},\;\;0\xrightarrow{\kappa_{4}}X_{1}.$ (6)
Of course, network (6) should be considered equivalent to network (1). So, it is also equivalent to network (2). Hence, in our context, we say that two networks are equivalent if they have the same steady-state ideal after undergoing relabeling, see Definition 3 (for further details on the “dynamical equivalence” of networks, refer to [5]). Determining equivalence among a large class of reaction networks is computationally challenging, as the standard approach involves computing their respective reduced Gröbner bases. In a recent related work [13], the authors point out that there are too few works investigating the equivalence of steady-state ideals, and we also see that deriving explicit conditions for the equivalence is difficult in general.
In this work, we propose focusing on a specific class of small networks to make this problem more feasible. The idea of studying small networks is motivated by the fact that many important dynamical features are inherited from a large network to a smaller “subnetwork” such as multistability [18, 7], oscillation [2] and local bifurcations [6]. A growing body of recent research focuses on finding the smallest networks from extensive classes of networks that admit these dynamical behaviors, and determining network equivalence first is essential to simplify the search process [3, 26, 19, 16]. Although computing the reduced Gröbner basis [1] for a single small network might not be difficult, determining equivalence in the aforementioned studies remains challenging, as one typically needs to handle a substantial amount of networks.
Another core proposal of this work is to focus specifically on zero-one networks for equivalence determination. Zero-one networks are crucial for intracellular signal transduction primarily such as the phosphorylation cycle [15], cell cycle [23], hybrid histidine kinase [25], and so on. They inherently ensure signal robustness, enabling stable responses against minor internal or external fluctuations instead of ambiguous intermediate outputs, which is essential for cells to make accurate functional decisions (e.g., proliferation, apoptosis).
In this work, our main contribution is an efficient algorithm (Algorithm 1 with sub-algorithms Algorithms 23) for determining the equivalence of zero-one networks that admit positive steady states. The efficiency of the proposed new method comes from several criteria for determining the equivalence of the steady-state ideals arising from zero-one networks, which dramatically cuts down the expenses on computing Gröbner bases since many computations are avoided. Note that zero-one networks are enumerable when the numbers of species and reactions are given. We make use of this fact to generate the possible rational vectors “aa” stated in Theorem 1, enabling efficient verification of whether a network admits positive steady states, see Algorithm 2. By a similar idea, we generate the possible rational vectors “aa” stated in condition (I) of Theorem 2 to efficiently check if a steady-state ideal is vacuous, see Algorithm 3. In some special cases, the “zero-one” assumption can be used to develop simple criteria for determining the vacuous ideals, see Theorem 3. Consequently, our proposed method obviates the necessity of Gröbner basis computations, and accurately excludes both networks that do not admit positive steady states and those with vacuous ideals, thus enabling highly efficient, precise pre-screening of target networks.
This paper reports improved and extended research work relative to the preliminary version published at ISSAC 25, with specific advancements as follows. First, in terms of algorithm design and implementation, we make the following two main improvements.
  • 1.
    We have refined the first three steps of the main algorithm [16, Algorithm 1]. We enumerate stoichiometric matrices instead of reaction networks, motivated by two key observations: a single stoichiometric matrix typically maps to numerous distinct networks, and preliminary steps including consistency verification are only necessary for stoichiometric matrices.
  • 2.
    We develop Theorem 1 and Algorithm 2 to verify consistency, given that most input networks do not admit positive steady states and we aim to filter them out prior to determining equivalence.
We have retested the new algorithm to determine the equivalence of the smallest zero-one networks that admit multistationarity (three-dimensional, three-species and five-reaction [17]). As reported in [16], this class contains over three millions networks, and the new algorithm returns $38185$ networks after determining the equivalence. Experimental results confirm the new algorithm achieves $50\%$ higher efficiency, halving computational time relative to the old version. A comparison between [16, Table 1] and Table 1 reveals a reduction in total computational time from $1$ hour to less than $0.5$ hours. It is seen from Table 1 that although the total number of tested networks exceeds three millions, the number of corresponding stoichiometric matrices is merely $43520$. After implementing the first three steps, the number is further reduced to $2548$ as the matrices that admit no positive steady states are excluded and the natural equivalence is checked. At this stage, the corresponding network count drops to around $100,000$. Notably, the first three steps take only $3$ minutes to complete, while the old version took roughly half an hour to perform these preliminary operations.
Second, from an application standpoint, we further apply our algorithm to two important families of small zero-one networks: three-dimensional systems (three species with six reactions) and four-dimensional systems (four species with five reactions). Our overarching goal is to identify all minimal zero-one networks that admit multistability [17] and Hopf bifurcations [26] from these two families, respectively. Given the substantial size of each family, our current testing efforts have focused exclusively on quadratic networks. Computational results show that no networks remain in either family after excluding inconsistent networks, those with vacuous steady-state ideals, and those with monomial steady-state ideals (see Tables 23). This implies that for generic choices of rate constants, such small quadratic zero-one networks lack positive steady states and thus cannot display nontrivial dynamical behaviors (e.g., multistability and periodic orbits). Notably, this finding aligns with the established result for two-dimensional quadratic networks [6, Lemma 3], suggesting that dynamically nontrivial small zero-one networks must be at least cubic.
@@
The rest of this paper is organized as follows. Section 2 reviews core concepts of reaction networks and the dynamical systems derived from mass-action kinetics, while formally defining the steady-state ideal and network equivalence within our framework. Section 3 establishes a suite of practical criteria—including Theorems 13 and Lemma 1—for determining network equivalence. Section 4 presents the main algorithm (Algorithm 1) alongside its auxiliary sub-algorithms (Algorithms 23) tailored to verifying the equivalence of zero-one networks. Section 5 demonstrates the main algorithm’s performance via applications to three-dimensional zero-one networks (three species, five reactions), as well as two families of quadratic zero-one networks: three-dimensional (three species, six reactions) and four-dimensional (four species, five reactions). This section further elaborates on implementation details and reports computational runtime results.
@@ 2 Background
where all stoichiometric coefficients $\alpha_{ij}$ and $\beta_{ij}$ are non-negative integers, and we assume that $(\alpha_{1j},\ldots,\alpha_{sj})\neq(\beta_{1j},\ldots,\beta_{sj})$. Each $\kappa_{j}\in\mathbb{R}_{>0}$ is called a rate constant corresponding to the $j$-th reaction in (7). We say a reaction is a zero-one reaction, if the stoichiometric coefficients $\alpha_{ij}$ and $\beta_{ij}$ in (7) belong to $\{0,1\}$. We say a network (7) is a zero-one network if it only has zero-one reactions. We say a reaction is quadratic, if the sum of stoichiometric coefficients $\alpha_{ij}$ in (7) equals to $2$. We say a network (7) is a quadratic network if it only has quadratic reactions. For any reaction defined in (7), the stoichiometric vector is $(\beta_{1j}-\alpha_{1j},\dots,\beta_{sj}-\alpha_{sj})^{\top}\in\mathbb{Q}^{s}$. We say finitely many reactions are linearly independent if their corresponding stoichiometric vectors are linearly independent. We call the $s\times m$ matrix with $(i,j)$-entry equal to $\beta_{ij}-\alpha_{ij}$ the stoichiometric matrix of $G$, denoted by $\mathcal{N}$. The dimension of $G$ is defined as the rank of $\mathcal{N}$. We call the $s\times m$ matrix the reactant matrix of $G$ with $(i,j)$-entry equal to $\alpha_{ij}$, denoted by $\mathcal{Y}$.
@@
$\dot{x}\;=\;f(\kappa,x)\;:=\;\mathcal{N}v(\kappa,x),$ (8)
@@
Treating the rate constants as a real vector of parameters $\kappa:=(\kappa_{1},\dots,\kappa_{m})^{\top}$ yields polynomials $f_{i}(\kappa,x)\in\mathbb{Q}(\kappa)[x]$, for $i\in\{1,\dots,s\}$. The polynomial ideal generated by $f$ in $\mathbb{Q}(\kappa)[x]$, denoted by $\langle f\rangle$, is called the steady-state ideal.
@@
For two $s\times m$ matrices $\mathcal{A}$ and $\hat{\mathcal{A}}$, we say the matrix $\hat{\mathcal{A}}$ has the form of the matrix $\mathcal{A}$ if we can obtain $\hat{\mathcal{A}}$ from $\mathcal{A}$ by permuting the rows and columns.
Definition 2.
For two networks $G$ and $\hat{G}$, we say the network $\hat{G}$ has the form of $G$ if we can obtain $\hat{G}$ from $G$ by relabeling the species $X_{1},\ldots,X_{s}$ as $\hat{X_{1}},\ldots,\hat{X_{s}}$, or by relabeling the reactions ${\mathcal{R}}_{1},\ldots,{\mathcal{R}}_{m}$ as $\hat{{\mathcal{R}}}_{1},\ldots,\hat{{\mathcal{R}}}_{m}$.
Definition 3.
Two networks $G_{1}$ and $G_{2}$ are equivalent if there exist $\hat{G_{1}}$ and $\hat{G_{2}}$ such that $\hat{G_{i}}$ has the form of $G_{i}$ for $i\in\{1,2\}$, and $\hat{G_{1}}$ and $\hat{G_{2}}$ have the same steady-state ideal.
For any given rate-constant vector $\kappa^{*}\in{\mathbb{R}}^{m}_{>0}$, a steady state of (8) is a vector of concentrations $x^{*}\in\mathbb{R}_{\geq 0}^{s}$ such that $f(\kappa^{*},x^{*})=\vec{0}$, where $f(\kappa,x)$ is on the right-hand side of the ODEs (8), and $\vec{0}$ denotes the column vector whose coordinates are all zero. If all the coordinates of a steady state $x^{*}$ are strictly positive (i.e., $x^{*}\in\mathbb{R}_{>0}^{s}$), then we call $x^{*}$ a positive steady state. In applications, we usually only care about positive steady states. We say a network admits positive steady states (or, consistent) if there exists a rate-constant vector such that system (8) has at least one positive steady state. For the stoichiometric matrix $\mathcal{N}\in\mathbb{Q}^{s\times m}$, the positive flux cone of $\mathcal{N}$ is defined as
@@
It is well-known that a network admits positive steady states if and only if $\mathcal{F}^{+}(\mathcal{N})\neq\emptyset$.
@@ 3 Theorem
3.1 Sufficient conditions for inconsistent
Theorem 1.
Consider a network $G$ with a stoichiometric matrix $\mathcal{N}$. If there exists $a\in\mathbb{Q}^{s}$ such that $a^{\top}\mathcal{N}$ is non-zero and does not change sign, then the network $G$ admits no positive steady states.
Proof.
Let $f$ be the steady-state system defined as in (8). Let $v=(v_{1},\dots,v_{m})^{\top}$ be defined as in (9). By (8), we have
$a^{\top}\mathcal{N}v(\kappa,x)=\sum\limits_{i=1}^{s}a_{i}f_{i}(\kappa,x),$ (11)
where $a_{i}$ denotes the coordinates of $a$. Assume that the network $G$ admits positive steady states. We prove the conclusion by deducing a contradiction. Note that $a^{\top}\mathcal{N}$ is non-zero and does not change sign. Hence, for any rate constant vector $\kappa\in\mathbb{Q}^{m}_{>0}$ and for any corresponding positive steady state $x\in\mathbb{Q}^{s}_{>0}$, the coefficients in $a^{\top}\mathcal{N}v(\kappa,x)$ are all positive or all negative. By (9), for any rate constant vector $\kappa\in\mathbb{Q}^{m}_{>0}$ and for any corresponding positive steady state $x\in\mathbb{Q}^{s}_{>0}$, we have $v(\kappa,x)\in\mathbb{Q}^{m}_{>0}$. Hence, $a^{\top}\mathcal{N}v(\kappa,x)$ is positive or negative. Then, by (11), for any rate constant vector $\kappa\in\mathbb{Q}^{m}_{>0}$ and for any positive steady state $x\in\mathbb{Q}^{s}_{>0}$, $\sum\limits_{i=1}^{s}a_{i}f_{i}(\kappa,x)$ is positive or negative. It is contrary to $x$ is a steady state. Hence, the network $G$ admits no positive steady states. ∎
Remark 1.
We remark that the hypotheses of Theorem 1 can be computationally checked by algebraic tools such as FindInstance in Mathematica [21]. Also note that if a matrix $\mathcal{N}$} has a non-zero row that does not change sign, then the matrix satisfies the hypotheses of Theorem 1. See the following example.
Example 1.
Consider the following zero-one network:
$0\xrightarrow{\kappa_{1}}X_{3},\;\;0\xrightarrow{\kappa_{2}}X_{2},\;\;0\xrightarrow{\kappa_{3}}X_{2}+X_{3},\;\;X_{1}+X_{2}+X_{3}\xrightarrow{\kappa_{4}}0.$
The stoichiometric matrix $\mathcal{N}$ is
$\begin{pmatrix}0&0&0&-1\\ 0&1&1&-1\\ 1&0&1&-1\end{pmatrix}$.
Let $a=(1,0,0)^{\top}$. Note that $a^{\top}\mathcal{N}=(0,0,0,-1)$. Hence, by Theorem 1, the network $G$ admits no positive steady states. On the other hand, note that $f_{1}=-\kappa_{4}x_{1}x_{2}x_{3}$, and so the steady-state system $f$ obviously has no positive solutions.